OpenStudy (anonymous):

x^-7x+9yi = y^i+20i-12 here ^ symbol means square find and y

OpenStudy (anonymous):

\[x ^{-7x+9yi} =y ^{i+20i-12}\] mean like this?

OpenStudy (anonymous):

noooo its not like thisssss

OpenStudy (anonymous):

see from here its q2 c part

OpenStudy (anonymous):

\[x ^{2}-7x+9yi=y ^{2}i+20i-12\]

OpenStudy (anonymous):

\[x ^{2}-7x+12=y ^{2}i+9yi+20i\] \[(x-3) (x-4) = i (y ^{2}+9y+20)\] \[(x-3) (x-4) = i (y+4) (y+5)\] Is this step help you to find x and y?

OpenStudy (anonymous):

wht will be the x and wht will be the y telll